OpenStudy (anonymous):

Can somebody please help me with Algebra 2 word problems? Any amount of help is greatly appreciated!

OpenStudy (anonymous):

I have a few problems that I can't figure out...

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

whats the question.?

OpenStudy (anonymous):

There are a few... Here's the first: Juice-O-Rama is a refreshment stand specializing in a fruit drink mixture of orange juice and prune juice. The orange juice costs $1.20 per liter and the prune juice costs $1.40 per liter. If they need 8 liters of the final drink at a cost of $1.24 per liter, how many liters of each juice will they need to make their specialty drink?

OpenStudy (anonymous):

I have no clue how to even START this... :(

OpenStudy (anonymous):

if they need 8 liters of 2 juices that makes 4 liters of each

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

1.20 x 4 = $ 1.24 x 4 = $ $ + $ = ?

OpenStudy (anonymous):

what's the $ for?

OpenStudy (anonymous):

? / 8 = amount per liter

OpenStudy (anonymous):

just used it rather than a ?

OpenStudy (anonymous):

I'm lost...

OpenStudy (anonymous):

1.20 is the amount per liter for orange juice and 1.24 is the amount per liter for orange juice

OpenStudy (anonymous):

right

OpenStudy (anonymous):

wait 1.40 is for prune

OpenStudy (anonymous):

so 1.2 is orange 1.4 is prune and 1.24 is the mix?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now if you were to get 8 liters of the mix you would need a certain amount of each juice to get $1.24 in profit from each liter of the mix

OpenStudy (anonymous):

yeah, but how do you figure that out?

OpenStudy (anonymous):

well i usually just go through all the possibilities and just plug them in to find the one i need 1.2a + 1.4p = x a + p = 8 x / 8 = 1.24

OpenStudy (anonymous):